Question7 of 10 A 0.00230 kg air-inflated balloon is given an excess negative charge q1 -3.75 x 10-8 C by rubbing it with a blanket. It is found that a charged rod can be held above the balloon at a distance of d = 0.0540 m to make the balloon float. In order for this to occur, what polarity of charge must the rod possess? O proton O neutral O negative positive How much charge q does the rod have? Model the balloon and rod as point charges. The Coulomb force constant is /)899x 10 Nm2/C2, and the acceleration due to gravity is g-9.81 m/s2
Answer #1

The charge has to be positive so that the attractive coulumbic force by the rod balances the weight of the balloon

kq1q2/r^2=mg

8.99*10^9*3.75*10^-8*q2/.054^2=.0023*9.81

q2=19.5*10^-8C